Cameroon Corp. manufactures and sells electric staplers for $17.00 each. If 10,000 units were sold in December, and management forecasts 5.0% growth in sales each month, the number of electric stapler sales budgeted for March should be:

Multiple Choice

  • 10,000

  • 11,576

  • 10,500

  • 11,025

  • 11,100
Answer #1

10,000 staplers are sold in December and sales increases every month by 5%

January sales = 10,000 + 5% = 10,500

February sales = 10,500 + 5% = 11,025

March sales = 11,025 + 5% = 11,576

2nd option